OpenStudy (anonymous):

Find the base two answer to 10101011two divided by 1001two?

OpenStudy (studygurl14):

Could you clarify that equation?

OpenStudy (anonymous):

Basically, "Find the quotient." \[\frac{10101011_2}{1001_2}=\cdots_2\]

OpenStudy (anonymous):

I forget the rules for base two arithmetic, so I'll just convert to base 10 first: \[\frac{10101011_2}{1001_2}=\frac{171_{10}}{9_{10}}\]

OpenStudy (anonymous):

which is \(19_{10}\). Should be easy enough to convert that into binary. I'm leaving that to you.

OpenStudy (amistre64):

10011 ------------ 1001 | 10101011 -1001 ------ 1101 -1001 ------ 1001 -1001 ------ 0